OpenStudy (anonymous):

2 cos(θ) + 9 = 8 equation on the interval 0< θ < 2π solving for θ

OpenStudy (anonymous):

\[2\cos(\theta)+9=8\\ 2\cos(\theta)=-1\\ \cos(\theta)=-\frac{1}{2}\] is a start then find the two values of \(\theta\) in \(0\leq \theta<2\pi\) where \(\cos(\theta)=-\frac{1}{2}\)

OpenStudy (anonymous):

so would it be 2π/3 and 4π/3